# On-the-fly training of polynomial machine learning potentials in computing lattice thermal conductivity

### v1

@@ IV.3 Comparison with conventional LTC calculation
In Fig. 5, the LTC values of the 103 compounds calculated through the polynomial MLPs trained using the 20 supercell datasets are compared with those calculated in the conventional approach using the same finite-difference displacement-force datasets Togo as those employed in Ref. Togo, 2023. These datasets share the same unit cells and supercell sizes for each compound. The latter datasets for the wurtzite-, zincblende-, and rocksalt-type compounds consist of 1254, 222, and 146 displacements, respectively, with a displacement distance of 0.03 Å. These displacements were systematically introduced considering crystal symmetries Chaput et al. (2011) by using the phono3py code.Togo et al. (2015, 2023) In all these calculations, the same version of the phono3py code Togo et al. (2015, 2023) was utilized to calculate the LTCs from the respective supercell force constants. The results demonstrate that the LTC values obtained through the polynomial MLPs consistently agree with those predicted by the conventional approach.Togo (2023)
@@
Figure 5: Comparison between LTC values calculated through the polynomial MLPs and those by the conventional approach.Togo (2023) To train the polynomial MLPs for each compound, we employed displacement-force pairs and energies of 20 supercells obtained through first-principles calculations. For the latter LTCs, we used the finite-difference displacement-force datasets Togo from Ref. Togo, 2023, comprising 1254, 222, and 146 supercells for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively, to fit third-order supercell force constants. The forces in these datasets were computed by first-principles calculations in equivalent calculation configurations with those employed in preparing datasets for training the polynomial MLPs.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP21K04632, JP22H01756, and JP19H05787.

### v3

@@ IV.3 Comparison with conventional LTC calculation
In Fig. 5, the LTC values of the 103 compounds calculated through the polynomial MLPs trained using the 20 supercell datasets are compared with those calculated in the conventional approach using the same finite-difference displacement-force datasets [35] as those employed in Ref. 15. These datasets share the same unit cells and supercell sizes for each compound. The latter datasets for the wurtzite-, zincblende-, and rocksalt-type compounds consist of 1254, 222, and 146 displacements, respectively, with a displacement distance of 0.03 Å. These displacements were systematically introduced considering crystal symmetries [36] by using the phono3py code.[9, 17] In all these calculations, the same version of the phono3py code [9, 17] (release v3.0.3) was utilized to calculate the LTCs from the respective supercell force constants. The results demonstrate that the LTC values obtained through the polynomial MLPs consistently agree with those predicted by the conventional approach.[15] The LTC values of the 103 compounds are also tabulated in Tables 3 and 4.
@@
Figure 5: Comparison between LTC values calculated through the polynomial MLPs and those by the conventional approach.[15] These values are tabulated in Tables 3 and 4. To train the polynomial MLPs for each compound, we employed displacement-force pairs and energies of 20 supercells obtained through first-principles calculations. For the latter LTCs, we used the finite-difference displacement-force datasets [35] from Ref. 15, comprising 1254, 222, and 146 supercells for the wurtzite-, zincblende-, and rocksalt-type compounds, respectively, to fit third-order supercell force constants.
Table 3: LTC values of zincblende- and wurtzite-type compounds at 300 K in W/$\cdots$ as shown in Fig. 5 calculated through the polynomial MLPs, z-$\kappa^{\text{pMLP}}$ and w-$\kappa^{\text{pMLP}}$, respectively, and those by the conventional approach, z-$\kappa^{\text{conv}}$ and z-$\kappa^{\text{conv}}$, respectively.
z-$\kappa^{\text{conv}}$ z-$\kappa^{\text{pMLP}}$ w-$\kappa^{\text{conv}}_{xx}$ w-$\kappa^{\text{conv}}_{zz}$ w-$\kappa^{\text{pMLP}}_{xx}$ w-$\kappa^{\text{pMLP}}_{zz}$
AgI 3.538 3.260 2.491 3.008 2.229 2.654
AlAs 79.84 79.88 68.18 66.24 62.99 61.21
AlN 232.0 233.6 269.5 243.7 271.4 244.4
AlP 79.18 79.21 71.84 65.52 71.17 65.54
AlSb 91.03 91.19 52.59 49.54 49.09 46.89
BAs 1614 1605 1535 1155 1541 1092
BeO 361.7 361.0 317.8 325.0 316.8 322.9
BeS 194.2 194.4 165.7 149.1 153.6 141.8
BeSe 447.3 443.3 345.7 333.7 342.2 327.8
BeTe 262.3 262.3 214.4 202.3 200.5 191.6
BN 1301 1290 1058 1012 1043 972.2
BP 489.5 488.3 498.5 359.9 468.0 338.1
CdS 23.67 23.52 19.81 19.99 19.40 19.71
CdSe 15.38 15.26 12.53 13.19 10.61 11.31
CdTe 8.523 8.358 6.284 7.215 6.043 6.812
CuBr 8.219 6.795 3.177 4.711 2.832 3.845
CuCl 2.056 1.681 1.452 2.483 1.383 2.224
CuH 21.44 21.16 6.248 5.724 6.202 5.949
CuI 10.68 8.844 8.773 9.872 7.777 8.821
GaAs 36.66 36.19 33.52 30.27 32.68 30.01
GaN 263.1 266.7 318.0 308.3 306.9 289.8
GaP 128.4 129.4 139.7 118.9 130.9 110.7
GaSb 37.74 37.30 28.5 21.90 25.25 20.13
InAs 25.47 25.35 24.23 23.28 22.92 22.08
InN 103.3 103.7 116.8 120.6 112.1 113.2
InP 83.65 85.97 75.73 72.38 70.61 63.27
InSb 15.21 14.96 11.76 11.84 10.96 11.08
MgTe 11.65 11.04 8.263 9.050 7.834 8.341
SiC 459.9 460.8 512.6 426.4 508.4 425.1
ZnO 64.27 63.47 51.38 59.91 50.80 57.65
ZnS 60.17 58.97 59.81 63.11 55.63 59.30
ZnSe 20.55 20.25 18.77 18.88 17.61 17.39
ZnTe 29.79 31.08 21.54 21.43 19.78 19.56
Table 4: LTC values of rocksalt-type compounds at 300 K in W/$\cdots$ shown in Fig. 5 calculated through the polynomial MLPs (r-$\kappa^{\text{pMLP}}$) and those by the conventional approach (r-$\kappa^{\text{conv}}$).
r-$\kappa^{\text{conv}}$ r-$\kappa^{\text{pMLP}}$
AgBr 0.5341 0.5363
AgCl 0.5053 0.5095
BaO 3.267 3.280
BaS 7.348 7.347
BaSe 11.48 11.50
BaTe 11.28 11.30
CaO 27.52 29.18
CaS 37.10 37.23
CaSe 19.56 19.70
CaTe 11.92 11.94
CdO 9.794 9.845
CsF 1.668 1.420
KBr 2.880 2.869
KCl 7.208 7.230
KF 7.648 7.670
KH 11.58 11.72
KI 2.087 2.098
LiBr 3.057 3.058
LiCl 4.972 4.995
LiF 16.50 16.53
LiH 40.86 41.25
LiI 2.375 2.397
MgO 66.35 66.35
NaBr 3.268 3.263
NaCl 8.030 8.070
NaF 24.39 24.47
NaH 17.62 17.90
NaI 2.409 2.389
PbS 2.130 2.129
PbSe 2.331 2.331
PbTe 3.163 3.175
RbBr 3.501 3.518
RbCl 2.542 2.553
RbF 3.330 3.357
RbH 6.145 6.141
RbI 1.967 1.978
SrO 11.98 12.03
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP21K04632, JP22H01756, JP19H05787 and 24K08021 Some of the calculations in this study were performed on the Numerical Materials Simulator at NIMS.